OpenStudy (anonymous):

Solve. 13-0.3(j-9)=0.9(11j-9)

OpenStudy (whpalmer4):

Use the distributive property to expand both sides. Then add and subtract quantities from both sides to get j alone on one side. If there is a coefficient for j that does not equal 1 at that point, divide both sides by the coefficient and you'll be left with j = <the answer>. If you don't remember what the distributive property is: \[a(b-c) = ab -ac \] Hint: don't forget about subtracting negative numbers! \[-a(b-c) = -ab -(-ac) = -ab + ac\]

OpenStudy (spareb665):

j=2.3333